Find the principal needed to get $1100 in 8 years at 12% compounded monthly.

Answer by mananth(16946)

Principal P = x
Amount= 1100
years=n 8
compounded 12 times a year
Rate = 12.00 =(0.12/12)=0.010
Amount = P*((n+r)/n)^n

1100 = x *( 1 + r/n )^ n


1100 = P ( 1 + 0.010 ) ^ 12 * 8

1100 = P( 1.01000 )^ 96

1100 /( 1.010 )^ 96 = P

P= $423.20